# Security Weaknesses of Copilot Generated Code in GitHub

### v2

@@ 3.3.2. Data Analysis
Step 5. We used well-known automated static analysis tools listed by OWASP (OWASP, [n. d.]) to scan the collected code snippets. Static analysis has been widely used to detect security issues in code, for its ability to analyze programs without execution (Dunlap et al., 2023). While dynamic analysis offers more precise insights by reasoning about program behavior, it suffers from limited coverage (Sui et al., 2020). Employing static analysis enables us to analyze program segments (smaller snippets) without requiring the entire program to be executed. Since different static analysis tools may use different algorithms and rules to detect security weaknesses, using multiple tools can increase our chances of discovering security issues in the code. To improve the coverage and accuracy of the results, we used two static analysis tools for security checks on each code snippet (i.e., CodeQL plus one dedicated tool for the specific language, Bandit for Python and ESLint for JavaScript).
We first used CodeQL to analyze the code in our dataset. The default query suite for the standard CodeQL query package is codeql-suites/<lang>-code-scanning.qls. Each package has several query suites in the codeql-suite directory. For example, the codeql/python-queries package contains the following query suites (GitHub, 2023):
@@
In this study, we scanned code snippets using the <language>-security-and-quality.qls test suite. These test suites check for multiple security properties and cover many CWEs. For example, the python-security-and-quality.qls test suite for Python provides 168 security checks, and the JavaScript test suite provides 203 security checks. As the query reports only provide the name and description of the security issues, we manually matched the results in the query reports with the corresponding CWE IDs. We then selected one additional static security analysis tool for files in each programming language we analyzed: Bandit for Python and ESLint for JavaScript. As explained in Section 3.2, we considered the code snippet from the Repository label to be the entire code file, while the code snippet from the Code label exists in the code file, with the exact number of lines unspecified at this stage. Step 6. We scanned code snippets from the Repository and Code labels, and we filtered the scan results before analyzing them. In this step, we adopted part of the strategies used by developers when they perform different tasks with static analysis tools: warning prioritization and determining whether a warning is a false positive or a true report (Do et al., 2020). We first performed an initial filtering of the results based on the priority of the warnings. Specifically we removed the repeated scan results that were reported by the two tools, then removed the scan results that were not security issues, such as recommendations, to get the initial scan results. Subsequently, to confirm that the security issues were actually caused by the Copilot-generated code, the first author manually checked the scan results. For the initial results obtained under the Repository label, we checked them one by one. Specifically, we determined whether the corresponding location of the code snippet indeed had a security issue based on the line number information provided by the scan results. For the initial results obtained under the Code label, we confirmed whether the security issues truly existed and verified if the security issues were caused by Copilot-generated code. Specifically, after scanning the code file, we pinpointed the code snippet within the file based on the line number of the security issue indicated in the scan results. We assessed whether it was generated by Copilot by checking the surrounding comments and determined whether a security issue existed in that particular context. We further analyzed the filtered scan results in Step 7, detailed in Section 4, according to the specific RQs.
We provide the complete dataset (including code snippets, full scan results, and filtered results) in our replication package (Fu et al., 2024).
4. Results
We present the results of three RQs formulated in Section 3.1 below. For each RQ, we first explain how we analyzed the collected code snippets to answer the RQ. We then answer each of our three RQs.
4.1. RQ1: How secure is the code generated by Copilot?
Approach. To answer this RQ, we collected 452 Python and JavaScript code snippets generated by Copilot from GitHub projects. We used two static analysis tools (CodeQL + another language-dedicated tool, Bandit for Python and ESLint for JavaScript) to scan and analyze the code snippets and then combine the results obtained from the two tools. The aim is to achieve better coverage of security issues. Therefore, as long as one of the tools detected the presence of a security issue, the code snippet was considered vulnerable.
We identified three types of warnings from CodeQL analysis:
@@ 4.1. RQ1: How secure is the code generated by Copilot?
Since our research primarily focused on security weaknesses, we only counted code snippets that had warnings and errors, and we ignored recommendations on code quality. After obtaining the initial scan results related to security issues, we manually checked the scan results. For the scan results from the Repository label, we marked them as 1 (security issue exists) and 0 (no security issue exists). For the scan results from Code label, we marked them as 1 (security issue exists) and 0 (no security issue exists or the code snippet with a security issue was not generated by Copilot). Figure 7 shows the outcomes of the manual verification on the initial security scan results under the Repository and Code labels.
When manually checking the scan results from the Code label, we also needed to identify whether the security issues obtained from the scan were from Copilot-generated code snippets based on the comment (prompt message) that appears before the method. We provide a working example of the filtration of the scan results in Figure 8. In Step 1, we first went to the corresponding file to locate the specific code snippet based on the start line numbers of the scan results. In Step 2, we located the code at Line 149 in aggregation.py. We found that this code does have a security issue and determined that it was generated by Copilot based on the prompt messages above. Consequently, we marked the corresponding security scan result as “1”. We also located the code at Line 145 in _lytools.py. We found that this code had a security issue, but there were no prompt messages that would indicate that it was generated by Copilot. Consequently, we marked the corresponding security scan result as “0” and discarded this scan result from further analysis.
Finally, we kept the results marked as 1 and aggregated the filtered results obtained using multiple analysis tools to calculate the number of code snippets with security issues detected.
(a) Results from the Repository label
(b) Results from the Code label
Figure 7. Manual verification of the initial security scan results for the Repository and Code label
@@
Figure 8. Example of filtering scan results from the Code label that are generated by Copilot
Results. Table 3 shows the numbers of code snippets for different types and the numbers and percentages of code snippets with security weaknesses. From the statistical results, we found that out of the 452 code snippets generated by Copilot, 25.9% of them have security weaknesses. Among the Python and JavaScript code snippets, Python exhibits a higher proportion of security issues compared to JavaScript. Out of the 277 Python code snippets, 91 (32.8%) have security weaknesses. Among the 175 JavaScript code snippets we collected, 43 (24.6%) have security weaknesses. Note that one related code snippet may contain multiple instances of a specific CWE.
Table 3. The number and percentage of code snippets with security weaknesses generated by Copilot
Language # Snippets
@@
Python 277 91 32.8%
JavaScript 175 43 24.6%
Total 452 134 29.6%
@@
4.2. RQ2: What security weaknesses are present in the code snippets generated by Copilot?
Approach. To answer RQ2, we processed the scan results collected by RQ1 that were manually checked to contain vulnerabilities. In total, we found 544 security issues in 452 code snippets, and Table 4 shows the number of security weaknesses found in code files of different programming languages.
Table 4. The number of security weaknesses in code snippets generated by Copilot
Language
# Snippets containing
security weaknesses
# Total security
weaknesses
Python 91 277
JavaScript 43 175
Total 134 544
@@ 4.2. RQ2: What security weaknesses are present in the code snippets generated by Copilot?
For each code snippet, we used CWEs to classify the security issues identified by the static analysis tools. Each CWE has a unique ID and a set of related descriptions, including its potential impact and how to detect and fix the CWE (The MITRE Corporation, 2023b). To more accurately identify the type of security issue with the corresponding CWE, we did not directly accept the provided information of the static analysis tools (such as the CWE number reported by Bandit). Instead, we manually identified the type of security issue at the specific location reported in the warning message of the static analysis tool and matched it with the corresponding CWE number. Initially, two authors independently matched each description of the security issue with a CWE ID. In case of disagreement, a discussion was initiated between the two authors, and one other author (a security expert) was then involved to provide his assessment. This process continued until all the descriptions of the security issues in the results were matched with CWE IDs. In the final stage, we performed a statistical analysis of CWE weaknesses in 134 code snippets that contained security weaknesses.
Results. Table LABEL:Distribution_of_CWEs_in_code_snippets shows the distribution of CWEs in the code snippets and the total number of occurrences (Frequency) of the CWE in the code snippets (we put those CWEs whose Frequency = 1 in “Others”). In total, we found 544 CWEs in 452 code snippets. These security weaknesses were related to 38 CWEs, indicating that developers face a variety of security weaknesses when using Copilot. CWE-330: Use of Insufficiently Random Values is the most frequently occurring CWE, as it represents 23.3% of the security weaknesses, followed by CWE-94: Improper Control of Generation of Code (‘Code Injection’) , CWE-78: OS Command Injection and CWE-95: Improper Neutralization of Directives in Dynamically Evaluated Code (‘Eval Injection’). Some CWEs appear less frequently, for example CWE-117: Improper Output Neutralization for Logs only occurred twice. Furthermore, many CWEs occur with a frequency of less than 1%, for example, CWE-22: Improper Limitation of a Pathname to a Restricted Directoryand CWE-770: Allocation of Resources Without Limits or Throttling. This indicates that the types of security issue are closely related to the specific scenarios in which developers use Copilot, emphasizing the importance of maintaining vigilance and caution when programming. Besides, Table 6 presents the top 5 CWEs that appear in Python and JavaScript code snippets.
@@
CWE-ID Frequency of Specific CWE Percentage
CWE-330 127 23.3%
CWE-94 115 21.1%
CWE-78 81 14.9%
CWE-95 30 5.5%
CWE-772 23 4.2%
CWE-89 15 2.8%
CWE-457 15 2.8%
CWE-259 14 2.6%
CWE-670 13 2.4%
CWE-396 11 2.0%
CWE-617 8 1.5%
CWE-628 7 1.3%
CWE-563 7 1.3%
CWE-561 7 1.3%
CWE-798 6 1.1%
CWE-703 6 1.1%
CWE-502 6 1.1%
CWE-20 6 1.1%
CWE-312 5 0.9%
CWE-605 4 0.7%
CWE-295 4 0.7%
CWE-252 4 0.7%
CWE-367 3 0.5%
CWE-327 3 0.5%
CWE-185 3 0.5%
CWE-79 2 0.4%
CWE-682 2 0.4%
CWE-665 2 0.4%
CWE-400 2 0.4%
CWE-215 2 0.4%
CWE-200 2 0.4%
CWE-117 2 0.4%
CWE-116 2 0.4%
Others =1 ¡1%
38 Types Total: 544
Table 6. Top 5 CWEs in Python and JavaScript

### v3

@@ 3.3.1. Data Analysis
Step 4. We used well-known automated static analysis tools suggested by OWASP (OWASP, 2024) to scan the collected code snippets. The reason why we chose a static rather than dynamic approach is that the static analysis has higher coverage and is able to analyze programs without executing them (Dunlap et al., 2023). While dynamic analysis offers more precise insights by reasoning about program behavior, it suffers from limited coverage and can be expensive (Sui et al., 2020; Ernst, 2003; Tahir and MacDonell, 2012). Static security analysis tools have been widely used by developers and companies (Distefano et al., 2019; Sadowski et al., 2018). Employing static analysis enables us to analyze program segments (smaller snippets) without requiring the entire program to be executed. Since different static analysis tools may use different algorithms and rules to detect security weaknesses, using multiple tools can increase our chances of discovering security weaknesses in the code. To improve the coverage and accuracy of the results, we used two static analysis tools for security checks on each code snippet (i.e., CodeQL plus one dedicated tool for the specific language, Bandit for Python, and ESLint for JavaScript).
CodeQL is a scalable static security analysis tool that is widely used in practice and allows users to analyze code and detect relevant weaknesses using predefined queries and test suites and supports for multiple languages (including Java, JavaScript, C++, C#, Go, and Python (GitHub, 2023)). Before using CodeQL to scan the identified code snippets for security weaknesses, we needed to create a CodeQL database for the source code. Source code can be directly analyzed for interpreted languages like Python and JavaScript without being compiled into intermediate code. Thus, we directly use the command line to generate the database needed for CodeQL queries. Then, we used CodeQL to analyze the generated database. We can find the available query suites for different programming languages in the codeql repository. For example, the codeql/python-queries package contains the following query suites (GitHub, 2023):
@@
In this study, we scanned code snippets using the <language>-security-and-quality.qls test suite. These test suites check for multiple security properties and cover many CWEs. For example, the python-security-and-quality.qls test suite for Python provides 168 security checks, and the JavaScript test suite provides 203 security checks. As the query reports only provide the name and description of the security weaknesses, we manually matched the results in the query reports with the corresponding CWE IDs based on the official documentation provided (CodeQL, 2024).
For Python files, we used Bandit (PyCQA, 2024) – a Python’s static security analysis tool. When running Bandit, we enabled all security rules by default. Bandit had a total of 73 security check rules, and each Bandit rule was typically associated with a corresponding CWE number, helping developers better understand the categories of detected vulnerabilities and align with industry standards. For JavaScript files, we used ESLint (ESLint Contributors, 2024) – a widely used JavaScript static analysis tool that can detect common coding issues and security vulnerabilities. Since we needed to configure the security check rules ourselves, we used the eslint-plugin-security plugin for security checks, which is an ESLint plugin containing 14 security check rules designed to detect common security vulnerabilities in JavaScript. In addition, we also configured the plugins eslint-plugin-no-secrets and eslint-plugin-xss with more precise rules to identify security weaknesses more accurately. For example, some of the rules in eslint-plugin-security also deal with XSS protection (such as avoiding unsafe eval() usage), but eslint-plugin-xss provides more precise rules against front-end XSS attacks. Both tools accept raw code snippets as input, and there is no need to pre-process the data before using them.
As explained in Section 3.2, we considered the code snippet from the Repository label to be the entire code file, while the code snippet from the Code label exists in the code file, with the exact location of the code snippets unspecified at this stage.
3.3.2. Results Filtering
Step 5. We scanned code snippets from the Repository and Code labels, and we filtered the analysis results before analyzing them. In this step, we adopted part of the strategies used by developers when they perform different tasks with static analysis tools: warning prioritization and determining whether a warning is a false positive or a true report (Do et al., 2020). We first performed an initial filtering of the results based on the priority of the warnings. Specifically, we manually removed the repeated analysis results that were reported by the two tools (14 duplicate results by Bandit and 2 duplicate results by ESLint), keeping only one for further analysis. The first author eliminated analysis results that were not security weaknesses. We identified three types of warnings from CodeQL analysis:
@@ 3.3.2. Results Filtering
Since our research primarily focused on security weaknesses, we only counted code snippets that had warnings and errors, and we ignored recommendations on code quality.
After obtaining preliminary analysis results related to security weaknesses, we manually checked the analysis results to confirm that the security weaknesses were actually caused by the automatically generated code. For the initial results obtained under the Repository label, we checked them one by one. Specifically, we determined whether the corresponding location of the code snippet indeed had a security weakness based on the line number information provided by the analysis results. For the initial results obtained under the Code label, we confirmed whether the security weaknesses truly existed and verified if the security weaknesses were caused by Copilot-generated code. Specifically, after scanning the code file, we pinpointed the code snippet within the file based on the line number of the security weakness indicated in the analysis results. We assessed whether it was generated by Copilot or other generation tools by checking the surrounding comments and determined whether a security weakness existed in that particular context. We further analyzed the filtered analysis results in Step 6, detailed in Section 4, according to the specific RQs.
3.4. Fixing Code with Security Weaknesses
Step 7. Based on RQ2 results, we randomly selected a subset of code snippets containing security weaknesses as the repair dataset since our purpose is to explore the feasibility of Copilot Chat in fixing the security weaknesses. These code snippets were all generated by Copilot, as most of the code obtained from GitHub was produced by Copilot. Since we aimed to fix as many security weaknesses as possible, we did not use clean data (i.e., code snippets without known security weaknesses). We randomly selected 50 Python and 40 JavaScript code snippets from the Repo label, which contained 163 and 132 security weaknesses, respectively. We utilized Copilot Chat to remediate the security weaknesses identified by the static analysis tools. We conducted several experiments on the selected dataset with 295 security weaknesses using three different prompts. Copilot Chat provides developers with basic slash commands, avoiding the need to write complex prompts for common scenarios. The slash commands include: /tests: generates unit tests for the selected code. /fix: provides fix suggestions for issues in the selected code. /explain: explains the selected code. /optimize: analyzes and improves the runtime of the selected code. We first used the slash command /fix to provide fix suggestions for the selected code. Then, we performed the fixes using our designed basic prompts and enhanced prompts with additional context. The prompt designs are as follows:
Basic Prompt Is there a security weakness with this code snippet? If so, please fix it and return the fixed code.
Enhanced Prompt This code snippet is known to have some security weaknesses. The warning messages are as follows: <insert warning messages here>. Please fix these security weaknesses and return the fixed code.
Figure 10. An example of using Copilot Chat to fix security weaknesses in Visual Studio Code
Figure 10 shows an example of how to use Copilot Chat to fix the security weaknesses in the code. We first select code snippets with security weaknesses and provide the three prompts (/fix, basic prompt, and enhanced prompt) in the IDE chat window. Then, Copilot Chat uses the selected code as a reference (input) to generate responses to our prompts, which we can accept or reject. Considering the token limitation issue, we always selected the function block containing the security weaknesses as the reference for Copilot Chat. After accepting the fix suggestions from Copilot Chat, we used the static analysis tools to recheck the modified code snippets. We recorded the positions of the modified code, indicating where the accepted code is located. Then we examined whether security weaknesses still exist in the corresponding positions of the modified code according to the rechecking results of the static analysis tools.
We provided the complete dataset (including code snippets, full analysis results, and filtered results) in our replication package (Fu et al., 2024).
4. Results
In this section, We present the results of the three RQs formulated in Section 3.1. For each RQ, we first explain how we analyzed the collected code snippets to answer the RQ, and we then answer each of our three RQs.
4.1. RQ1: How secure is the code generated by Copilot in GitHub projects?
Approach. To answer this RQ, we collected 733 Python and JavaScript code snippets generated by Copilot and two other tools from GitHub projects, with 672 from Copilot, 38 from CodeWhisperer, and 23 from Codeium. We used two static analysis tools (CodeQL + another language-dedicated tool, Bandit for Python, and ESLint for JavaScript) to scan. The aim is to achieve better coverage of security weaknesses. Therefore, as long as one of the tools detected the presence of a security weakness, the code snippet was considered vulnerable.
Subsequently, we filtered the analysis results as described in Section 3.3.2. For the analysis results from the Repository label, we marked them as 1 (security weakness exists) and 0 (no security weakness exists). For the analysis results from Code label, we marked them as 1 (security weakness exists) and 0 (no security weakness exists or the code snippet with a security weakness was not generated by code generation tools).
@@ 4.1. RQ1: How secure is the code generated by Copilot in GitHub projects?
Figure 11. Example of filtering analysis results from the Code label that are generated by Copilot
When manually checking the analysis results from the Code label, we also needed to identify whether the security weaknesses obtained from the scan were from automatically generated code snippets based on the comment (prompt message) that appears before the method. We provide a working example of the filtration of the analysis results in Figure 11. In Step 1, we first went to the corresponding file to locate the specific code snippet based on the start line numbers of the analysis results. In Step 2, we located the code at Line 149 in aggregation.py. We found that this code does have a security weakness and determined that it was generated by Copilot based on the prompt messages above. Consequently, we marked the corresponding security analysis result as “1”. We also located the code at Line 145 in _lytools.py. We found that this code had a security weakness, but there were no prompt messages that would indicate that it was generated by Copilot. Consequently, we marked the corresponding security analysis result as “0” and discarded this analysis result from further analysis.
Two authors independently filtered the analysis results within a period of two weeks, and if there were any results that they were unsure of or disagreed with, the two authors discussed the results until they reached an agreement. Cohen’s Kappa coefficients (Cohen, 1960) were 0.85 for the analysis results from the Repository label, and 0.82 for the analysis results from the Code label, which were both higher than 0.8, ensuring a good accuracy of the filtering results. Finally, we kept the results marked as 1 and aggregated the filtered results obtained using multiple analysis tools to calculate the number of code snippets with security weaknesses detected. All labeling results can be found in our online replication package (Fu et al., 2024).
Results. Table 4 shows the number of code snippets containing security weaknesses and the number of security weaknesses identified in the code snippets. From the statistical results, we found that out of the 733 generated code snippets, 27.3% of them contained security weaknesses. Around 30% of Python code snippets (124 of the 419) and around 25% of JavaScript code snippets (76 of the 314) contained security weaknesses. At the same time, we can see that more than half of the code snippets containing security weaknesses (102 of 200) have more than one security issue. In Figure 12, we present the distribution of application domains for the 200 code snippets (27.3% of all) identified with security weaknesses. We also found 628 security weaknesses in the 200 code snippets. Note that one code snippet may contain multiple security weaknesses attributed to multiple CWEs. Figure 13 shows the distribution of security weaknesses per code snippet. Our analysis reveals an average incidence of 3 security weaknesses per code snippet, irrespective of the programming language. The maximum count of security weaknesses within a single code snippet, which consisted of 1,595 LoC, reached 22.
Table 4. Statistics of the code snippets with security weaknesses
Language # Snippets # Security weaknesses
@@
# containing more than
one security weakness
Python 419 387 124 (29.5%) 65 (15.5%)
JavaScript 314 241 76 (24.2%) 37 (11.8%)
Total 733 628 200 (27.3%) 102 (13.9%)
(a) application domains of Python code
(b) application domains of JavaScript code
Figure 12. Application domains of the code snippets with security weaknesses
Figure 13. Distribution of the number of security weaknesses per code snippet
@@
4.2. RQ2: What security weaknesses are present in the code generated by Copilot across different application domains?
@@
Approach. To answer RQ2, we processed the analysis results collected by RQ1, which were manually checked to see if they contained security weaknesses. For each code snippet, we used CWEs to classify the security weaknesses identified by the static analysis tools. Each CWE has a unique ID and a set of related descriptions, including its potential impact and how to detect and fix the CWE (MITRE, 2024). To accurately identify the type of security weakness with the corresponding CWE, we did not directly accept the information provided by the static analysis tools (for example, only Bandit provided the CWE IDs of the security issues identified, and we did not follow the CWE IDs of 55 out of 235 security issues reported by Bandit). Instead, we manually identified the type of security weakness at the specific location reported in the warning message of the static analysis tool. We utilized the warning information provided by static analysis tools, such as the test_name and message fields indicated by CodeQL, in conjunction with relevant descriptions from the CWE, to perform manual correspondences. Initially, two authors independently matched each description of the security issue with a CWE ID within a period of ten days. Cohen’s Kappa coefficient (Cohen, 1960) was 0.82, which was higher than 0.8 and indicated a high level of agreement between the two authors. In case of disagreement, a discussion was initiated between the two authors, and one other author (a security expert) was then involved to provide his assessment. We provided correspondences between the warning prompt messages and CWEs in our online replication package (Fu et al., 2024). In the final stage, we performed a statistical analysis of CWE weaknesses in 200 code snippets that contained security weaknesses. Additionally, we analyzed the distribution of different CWEs across different application domains of the code snippets.
We also discussed whether the security weaknesses present in the code generated by Copilot are widely prevalent in the code produced during software development. The code snippets in our collected dataset were mainly generated in 2022 and 2023. To compare whether the security weaknesses in in our dataset are widespread in this period, we chose MITRE 2023 CWE Top-25 list (MITRE, 2023) as our baseline. We compared the CWEs obtained in RQ2 with the 2023 CWE Top 25.
@@
CWE-ID Name Frequency Percentage
CWE-330 Use of Insufficiently Random Values Weakness 114 18.15%
CWE-94 Improper Control of Generation of Code (’Code Injection’) 62 9.87%
CWE-79 Improper Neutralization of Input During Web Page Generation (’Cross-site Scripting’) 60 9.55%
CWE-78 Improper Neutralization of Special Elements used in an OS Command (’OS Command Injection’) 39 6.21%
CWE-427 Uncontrolled Search Path Element 35 5.57%
CWE-457 Use of Uninitialized Variable 30 4.78%
CWE-22 Improper Limitation of a Pathname to a Restricted Directory (’Path Traversal’) 29 4.62%
CWE-772 Missing Release of Resource after Effective Lifetime 29 4.62%
CWE-89 Improper Neutralization of Special Elements used in an SQL Command (’SQL Injection’) 27 4.30%
CWE-259 Use of Hard-coded Password 16 2.55%
CWE-685 Function Call with Incorrect Number of Arguments 14 2.23%
CWE-284 Improper Access Control 13 2.07%
CWE-312 Cleartext Storage of Sensitive Information 12 1.91%
CWE-390 Detection of Error Condition Without Action 11 1.75%
CWE-456 Missing Initialization of Variable 11 1.75%
CWE-607 Use of Wrong Operator in String Comparison 11 1.75%
CWE-770 Allocation of Resources Without Limits or Throttling 11 1.75%
CWE-20 Improper Input Validation 8 1.27%
CWE-665 Improper Initialization 8 1.27%
CWE-117 Improper Output Neutralization for Logs 7 1.11%
CWE-400 Uncontrolled Resource Consumption 7 1.11%
CWE-561 Dead Code 7 1.11%
CWE-732 Incorrect Permission Assignment for Critical Resource 6 0.96%
CWE-798 Use of Hard-coded Credentials 6 0.96%
CWE-215 Information Exposure Through Debug Information 5 0.80%
CWE-290 Authentication Bypass by Spoofing 5 0.80%
CWE-295 Improper Certificate Validation 5 0.80%
CWE-209 Information Exposure Through an Error Message 4 0.64%
CWE-252 Unchecked Return Value 4 0.64%
CWE-571 Expression is Always True 4 0.64%
CWE-605 Multiple Binds to the Same Port 4 0.64%
CWE-200 Information Exposure 3 0.48%
CWE-327 Use of a Broken or Risky Cryptographic Algorithm 3 0.48%
CWE-367 Time-of-check Time-of-use Race Condition 3 0.48%
CWE-563 Assignment of a Fixed Address to a Pointer 3 0.48%
CWE-116 Improper Encoding or Escaping of Output 2 0.32%
CWE-480 Use of Incorrect Operator 2 0.32%
CWE-502 Deserialization of Untrusted Data 2 0.32%
CWE-601 URL Redirection to Untrusted Site (’Open Redirect’) 2 0.32%
CWE-208 Observable Timing Discrepancy 1 0.16%
CWE-570 Expression is Always False 1 0.16%
CWE-628 Function Not Implemented Correctly 1 0.16%
CWE-682 Incorrect Calculation 1 0.16%
43 Types Total: 628
Results. Table LABEL:Distribution_of_CWEs_in_code_snippets shows the distribution of CWEs in the code snippets and the total number of occurrences (frequency) of the CWEs in the code snippets. In total, we found 628 CWEs in 200 code snippets with security weaknesses. These security weaknesses were related to 43 CWE types, indicating that developers face a variety of security weaknesses when using Copilot and other code generation tools. CWE-330: Use of Insufficiently Random Values is the most frequently occurring CWE, as it represents 18.15% of the security weaknesses, followed by CWE-94: Code Injection , CWE-79: Cross-site Scripting, and CWE-78: OS Command Injection. Some CWEs appear less frequently. For example, CWE-117: Improper Output Neutralization for Logs only occurred twice. Furthermore, many CWEs occur with a frequency of less than 1%, for example, CWE-117: Improper Output Neutralization for Logsand CWE-290: Authentication Bypass by Spoofing. This indicates that the types of security weaknesses are closely related to the specific scenarios in which developers use code generation tools, emphasizing the importance of maintaining vigilance and caution when programming. In addition, we highlighted those CWEs that are the “Top-25 CWEs” in bold. It is worth noting that 233 security weaknesses in the code snippet correspond to these eight CWEs, which belong to the Top-25 CWE list, indicating that the CWE Top-25 weaknesses are also prevalent in the code generated by Copilot and other tools. At the same time, we can see that CWE-79: Cross-site Scripting and CWE-78: OS Command Injection are among the most frequently occurring security weaknesses in our results and rank high on the Top-25 CWE list.
Table LABEL:tab:cwe_types presents the top 5 CWEs that appear in Python and JavaScript code snippets. We found that the detected CWEs vary across programming languages. CWEs in Python are mainly related to data processing and system calls, such as CWE-78: OS Command Injection and CWE-427: Uncontrolled Search Path Element. In contrast, CWEs in JavaScript are primarily associated with dynamic code generation problems and security issues in Web development, such as CWE-94: Code Injection and CWE-79: Cross-site Scripting.
Furthermore, Figure LABEL:fig:cwe-py and Figure LABEL:fig:cwe-js show the distribution of CWEs across different application domains. We can observe that the number and type of instances of CWE vary from one application domain to another. In addition to this, the distribution of CWEs within the same application domain varies by programming language. For example, in Python, the application domain with the most CWEs is Utility Tool, while in JavaScript, Web Applications have the most CWEs. Meanwhile, we can see that in the Web Application category, CWE-79: Cross-site Scripting is the most frequent in JavaScript code snippets, while CWE-89: SQL Injection is the most present in Python code snippets.